# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v3

@@
Linear structural vector autoregressive models can be identified statistically without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian. We show that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. We also discuss the problem of labelling the shocks, estimation of the parameters, and stationarity the model. The introduced methods are implemented to the accompanying R package sstvars. Our empirical application studies the effects of the climate policy uncertainty shock on the U.S. macroeconomy. In a structural logistic smooth transition vector autoregressive model consisting of two regimes, we find that a positive climate policy uncertainty shock decreases production and increases inflation in times of both low and high economic policy uncertainty, but its inflationary effects are stronger in the periods of high economic policy uncertainty.
@@ 1 Introduction
Smooth transition vector autoregressive (STVAR) models (Anderson and Vahid, 1998) have been widely adopted in empirical macroeconomics due to their ability to flexibly capture nonlinear data generating dynamics and gradual shifts in parameter values. Such variation may arise due to wars, business cycle fluctuations, or policy shifts, for example. Structural STVAR models are particularly useful, as they facilitate tracing out the causal effects of economic shocks, which may vary in time depending on the initial state of the economy as well as on the sign and size of the shock. In order to estimate their effects, the economic shocks need to be identified. Conventional identification methods typically rely on restrictive assumptions such as zero contemporaneous interactions among some of the variables or long-term neutrality of the shocks. Such restrictions with a clear interpretation are advantageous when based on economic reasoning, but often they are economically implausible and imposed just to achieve sufficient identification. To overcome the latter issue, statistical identification methods relying on the statistical properties of the data can be used.
@@
There are two main branches in the statistical identification literature: identification by heteroscedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others). Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions. However, identification by heteroscedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see, e.g., Lütkepohl and Netšunajev, 2017). This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses. To the best our knowledge, identifying nonlinear SVAR models by non-Gaussianity has not been studied in the previous literature.
@@
This paper contributes to the literature on identification by non-Gaussianity by extending the framework proposed by Lanne et al. (2017) to STVAR models. We also consider threshold vector autoregressive models of Tsay (1998) as special cases of STVAR models. In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally be time-varying to allow for time-variation in the impact responses of the variables to the shocks, which complicates identification. Nonetheless, similarly to Lanne et al. (2017), it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian. We show that under this condition, when the impact matrix of a STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be generalized to other suitable weight functions as well. In line with statistical identification literature, however, external information is required to labelling the identified structural shocks as economic shocks. Since the identification can, in our experience, be also weak with respect to the ordering and signs of the columns of the regime-specific impact matrices, we recommend adopting a blended identification strategy that combines identification by non-Gaussianity with additional information (cf. Carriero et al., 2024).
@@
Developments related to ours include the new framework of Morioka et al. (2021) called independent innovation analysis (IIA). The IIA estimates the (conditionally) contemporaneously independent innovations of a nonlinear SVAR model without imposing any specific functional form. Under fairly general assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations. While the IIA does seem promising, it is not obvious how to apply it in empirical macroeconometrics, as identification only up to such transformations is, in our understanding, insufficient for structural analysis. Moreover, Morioka et al. (2021) assume innovations follow a distribution in the exponential family, thereby excluding various alternatives. In contrast, our approach facilitates estimation of the parameters of an identified nonlinear SVAR model, enabling structural analysis without restricting the distribution of independent shocks beyond ruling out Gaussian and infinite-variance cases.
@@
Our empirical application studies the macroeconomic effects of climate policy uncertainty shocks and consider monthly U.S. data from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to CPU. We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU). Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the EPU index (Baker et al., 2016) as the switching variable. We find that a positive CPU shock decreases production and increases inflation in times both low and high EPU, but its inflationary effects are stronger in the periods of high EPU. Our results are, hence, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
@@
The rest of this paper is organized as follows. Section 2 presents the framework of reduced form STVAR models and provides examples of the covered models. Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and discusses the problem of labelling the shocks. Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood -based estimator using a three-phase procedure. Section 5 presents the empirical application and Section 6 concludes. Appendices provide detailed discussion on ergodic stationarity of the model; proofs for the stated lemmas, propositions, and theorem; details related to the proposed estimation procedure; a Monte Carlo study assessing the performance of the proposed estimator; and details related to the empirical application. Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
@@ 2 Reduced form STVAR models
$\mu_{m,t}$ $=\phi_{m,0}+\sum_{i=1}^{p}A_{m,i}y_{t-i},\quad m=1,...,M,$ (2.2)
where $\phi_{1,0},...,\phi_{M,0}\in\mathbb{R}^{d}$, $m=1,...,M$, are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on. The transition weights $\alpha_{m,t}$ are assumed to be either exogenous (nonrandom) or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. Thus, the types of transition weights typically used in macroeconomic applications are accommodated. The transition weights express the proportions of the regimes the process is on at each point of time and determine how the process shifts between them.
@@
Our framework of STVAR models accommodates various types of transition weights employed in empirical macroeconomics. To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
@@
In addition to specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2). For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constants across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m,0}=\phi_{0}$ for all $i=1,...,p$ and $M=1,...,M$. The model then simplifies to
$y_{t}=\phi_{0}+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.5)
The linear model in (2.5) is typically substantially easier to estimate and has much fewer parameters than the fully nonlinear specification in (2.1)-(2.2), so it is more feasible in large scale VAR analyses, but it can still accommodate shifts in the volatility regime through the error process $u_{t}$.
@@ 3.1 Structural STVAR models
Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al. (2017), our identifying restriction is that the shocks are mutually independent and that at most one of them is Gaussian. Under this assumption, unlike in the Gaussian case, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ generally lead to observationally distinct models, effectively identifying a static impact matrix and the shocks (Lanne et al., 2017). We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with $B_{y,t}$.
In order to parametrize the model directly with the impact matrix, we make the identity $u_{t}=B_{y,t}e_{t}$ explicit in Equation (2.1) as
$y_{t}=\sum_{m=1}^{M}\alpha_{m,t}\left(\phi_{m,0}+\sum_{i=1}^{p}A_{m,i}y_{t-i}\right)+B_{y,t}e_{t},\quad e_{t}\sim IID(0,I_{d},\nu),$ (3.2)
@@
The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero. This implies that invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$). To further improve intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero. In practice, the parameter space should be simply restricted in the estimation so that parameter values leading to $B_{y,t}$ that is singular for some $t$ are excluded.
@@
where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$. The conditional covariance matrix of $y_{t}$ can be thus described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$. If the process is completely in one of the regimes, i.e., $\alpha_{\tilde{m},t}=1$ for some $\tilde{m}\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{\tilde{m}}$, implying that this constitutes as the conditional covariance matrix of Regime $\tilde{m}$. When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends both on the regime-specific covariance matrices and the cross terms. Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ is a weighted sum of the covariance matrices of the regimes. Nonetheless, our model specification facilitates exploiting non-Gaussianity of the shocks in their identification, as we show in the next section.
@@ 3.2 Identification of the shocks by non-Gaussianity
The key identification assumption is that the structural shocks $e_{t}=(e_{1t},...,e_{dt})$ are a mutually independent and at most one of them is Gaussian. We also assume that $e_{t}$ is an IID sequence with zero mean and identity covariance matrix as is formally stated in the following assumption.
@@
The next lemma then shows that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns. This means that only the impact matrices that are obtained by reordering or changing the signs of the columns of $B_{y,t}$ are observationally equivalent. This lemma, which is partly similar to Proposition 1 in Lanne et al. (2017), is used in the following sections to show that the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified. Moreover, Lemma 2 also shows that if the transition weights are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic (2.4) process, then also the autoregressive (AR) as well as the weight function parameters are identified. Specific weight functions are assumed here for simplicity, but the result can be extended to other suitable weight functions as well.
@@
Moreover, suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$. If the following conditions hold:
@@
    the transition weights $\alpha_{m,t}$, $m=1,...,M$, are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic process (2.4) (with $M=2$ in the latter), and
@@
    $(\phi_{m,0},\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))\neq(\phi_{n,0},\text{vec}(A_{n,1}),...,\text{vec}(A_{n,p}))$ for all $m\neq n\in\{1,...,M\}$,
the parameters $\phi_{m,0},A_{m,1},...A_{m,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. For threshold models (2.3), also the parameters $r_{1},...,r_{m-1}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. For logistic models (2.4), the parameters $c,\gamma$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, where $z_{t}$ is the switching variable defined in (2.4).
@@
We have concluded that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1. In other words, at each $t$‚ changing the ordering or signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to an observationally distinct model. The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1. Identification of the AR and weight function parameters, in turn, is achieved only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, because in some special cases of degenerate data (e.g., if the regime does not switch), there is insufficient variation to distinguish the parameters.
If the impact matrix is time-invariant as in Lanne et al. (2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows from Lemma 2 that the structural shocks are identified up to ordering and signs. Such statistically identified shocks can then often be labelled as economic shocks based on external information (see Lanne et al., 2017). However, when the impact matrix varies in time, two complications arise. First, the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, and it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified. Second, since the identification of $B_{y,t}$ at each $t$ is only up to ordering and signs of its columns, its unique identification requires that the parameters in the functional form of $B_{y,t}$ are restricted to fix the ordering and signs of the columns of $B_{y,t}$.
To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$. To fix the ordering and signs of the columns of $B_{y,t}$, it suffices to fix the ordering and signs of the columns of $B_{1},...,B_{M}$. This requires such constraints to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{m}$ are not observationally equivalent at some $t$.
It is not immediately obvious what is the best strategy for fixing the ordering and signs of the columns of $B_{1},...,B_{M}$, as depending on the transition weights and distributions of the shocks, constraints that are not restrictive in some specifications can be overidentifying in others. Also, constraints that are identifying in some specifications may not be sufficient for identification in others. Therefore, it is useful to divide the discussion on identification into the cases of binary (i.e., $\alpha_{m,t}\in\{0,1\}$) and more general transition weights.

### v4

@@
Linear structural vector autoregressive models can be identified statistically without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian. We show that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. We also discuss the problem of labelling the shocks, estimation of the parameters, and stationarity the model. The introduced methods are implemented to the accompanying R package sstvars. Our empirical application studies the effects of the climate policy uncertainty shock on the U.S. macroeconomy. In a structural logistic smooth transition vector autoregressive model with two regimes, we find that a positive climate policy uncertainty shock decreases production and increases inflation in times of both low and high economic policy uncertainty, but its inflationary effects are stronger in the periods of high economic policy uncertainty.
@@ 1 Introduction
Linear vector autoregressive (VAR) models are a standard tool in empirical macroeconomics due to their ability to effectively capture constant dynamic relationships of the included variables. Macroeconomic systems may, however, exhibit variation in their dynamics, induced by wars, crises, policy shifts, or business cycle fluctuations, for example. Such structural changes cannot be accommodated by linear VAR models, and therefore, nonlinear VAR models that are able capture such dynamics are often employed. Nonlinear structural VAR (SVAR) models are particularly useful, as they facilitate tracing out the causal effects of economic shocks, which may vary in time depending on the initial state of the economy as well as on the sign and size of the shock (Kilian and Lütkepohl, 2017, Chapter 18). To estimate these effects, the economic shocks need to be identified. Conventional identification methods typically employed in nonlinear SVAR models rely on restrictive assumptions, such as zero contemporaneous interactions among some of the variables (see Kilian and Lütkepohl, 2017, Section 18.2.2). Such restrictions with a clear interpretation are advantageous when based on economic reasoning, but often they are economically implausible and imposed just to achieve sufficient identification. To overcome the latter issue, statistical identification methods relying on the statistical properties of the data can be used.
@@
There are two main branches in the statistical identification literature: identification by heteroscedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others). To the best our knowledge, this paper is the first to study identification by non-Gaussianity in nonlinear SVAR models. Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions. However, identification by heteroscedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see, e.g., Lütkepohl and Netšunajev, 2017). This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses.
@@
This paper contributes to the literature on identification by non-Gaussianity by extending the framework proposed by Lanne et al. (2017) to smooth transition vector autoregressive (STVAR) models (e.g., Anderson and Vahid, 1998), which is major class of nonlinear SVAR models. We also consider the threshold vector autoregressive model of Tsay (1998) as a special case of the STVAR model. The STVAR model can flexibly capture nonlinear data generating dynamics by accommodating multiple regimes and gradual shifts between them, governed by the transition weighs. In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally allow for time-variation in the impact responses of the variables to the shocks, which complicates identification. Nonetheless, similarly to Lanne et al. (2017), it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian. We show that under this condition, when the impact matrix of the STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be extended to other suitable weight functions as well.
In line with the statistical identification literature, external information is required to label the identified structural shocks as economic shocks. Our nonlinear setup has the additional complication that the same shock should be assigned to the same column of the impact matrix across all regimes. Therefore, since our experience shows that the identification can be weak with respect to the ordering and signs of the columns of the regime-specific impact matrices, we recommend adopting a blended identification strategy that combines identification by non-Gaussianity with additional identifying information (cf. Carriero et al., 2024).
@@
Developments related to ours include the new framework of Morioka et al. (2021) called independent innovation analysis (IIA). The IIA estimates the (conditionally) contemporaneously independent innovations of a nonlinear SVAR model without imposing any specific functional form. Under fairly general assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations. While the IIA does seem promising, it is not obvious how to apply it in empirical macroeconometrics, as identification only up to such transformations is, in our understanding, insufficient for structural analysis. Moreover, Morioka et al. (2021) assume the innovations follow a distribution in the exponential family, thereby excluding various interesting alternatives. In contrast, our approach facilitates estimation of the parameters of an identified nonlinear SVAR model, enabling structural analysis without restricting the distribution of independent shocks beyond ruling out Gaussian and infinite-variance cases.
@@
Our empirical application studies the macroeconomic effects of climate policy uncertainty (shocks) and considers monthly U.S. data from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to CPU. We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU). Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the EPU index (Baker et al., 2016) as the switching variable. We find that a positive CPU shock decreases production and increases inflation in times of both low and high EPU, but its inflationary effects are stronger in the periods of high EPU. Our results are, hence, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
@@
The rest of this paper is organized as follows. Section 2 presents the framework of reduced form STVAR models and provides examples of the covered models. Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and discusses the problem of labelling the shocks. Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood -based estimator using a three-step procedure. Section 5 presents the empirical application and Section 6 concludes. Appendices provide detailed discussion on ergodic stationarity of the model; proofs for the stated lemmas, propositions, and theorem; details related to the proposed estimation procedure; a Monte Carlo study assessing the performance of the proposed estimator; and details related to the empirical application. Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
@@ 2 Reduced form STVAR models
$\mu_{m,t}$ $=\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-i},\quad m=1,...,M,$ (2.2)
where $\phi_{1},...,\phi_{M}\in\mathbb{R}^{d}$ are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on. The transition weights $\alpha_{m,t}$ are assumed to be either exogenous (nonrandom) or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. Thus, the types of transition weights typically used in macroeconomic applications, including exogenous, logistic, and threshold weights discussed below, are accommodated. The transition weights express the proportions of the regimes the process is on at each point of time and determine how the process shifts between them.
@@
Our framework accommodates various types of transition weights employed in empirical macroeconomics. To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
@@
In addition to specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2). For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constant across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m}=\phi$ for all $i=1,...,p$ and $M=1,...,M$. The model then simplifies to
$y_{t}=\phi+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.5)
The linear model in (2.5) is typically substantially easier to estimate and has much fewer parameters than the fully nonlinear specification in (2.1)-(2.2), making it more feasible for large scale SVAR analyses, but it can still accommodate shifts in the volatility regime through the error process $u_{t}$.
@@ 3.1 Structural STVAR models
Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al. (2017), our identifying restriction is that the shocks are mutually independent and at most one of them is Gaussian. Under this assumption, unlike in the Gaussian case, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ generally lead to observationally distinct models, effectively identifying a static impact matrix and the shocks (Lanne et al., 2017). We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with $B_{y,t}$. Hence, we make the identity $u_{t}=B_{y,t}e_{t}$ explicit in Equation (2.1) as
$y_{t}=\sum_{m=1}^{M}\alpha_{m,t}\left(\phi_{m}+\sum_{i=1}^{p}A_{m,i}y_{t-i}\right)+B_{y,t}e_{t},\quad e_{t}\sim IID(0,I_{d},\nu),$ (3.2)
@@
The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero. This implies that invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$). To further improve intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero. In practice, the parameter space should be simply restricted in estimation so that parameter values leading to $B_{y,t}$ that is singular for some $t$ are excluded.
@@
where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$. The conditional covariance matrix of $y_{t}$ can be thus described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$. If the process is completely in one of the regimes, i.e., $\alpha_{m,t}=1$ for some $m\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{m}$, implying that this constitutes as the conditional covariance matrix of Regime $m$. When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends on both the regime-specific covariance matrices and the cross terms. Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ is a weighted sum of the covariance matrices of the regimes. Nonetheless, our model specification facilitates exploiting non-Gaussianity of the shocks in their identification, as we show in the next section.
@@ 3.2 Identification of the shocks by non-Gaussianity
The key identification assumption is that the structural shocks $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one of them is Gaussian. We also assume that $e_{t}$ is an IID sequence with zero mean and identity covariance matrix as is formally stated in the following assumption.
@@
The next lemma then shows that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns. This lemma, which is partly similar to Proposition 1 in Lanne et al. (2017), is used in the following sections to show that the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified. Moreover, Lemma 2 also shows that if the transition weights are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic (2.4) process, then also the autoregressive (AR) and weight function parameters are identified. Specific weight functions are assumed here for simplicity, but the result can be extended to other suitable weight functions as well.
@@
Moreover, suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$ and that the following conditions hold:
@@
    the transition weights $\alpha_{m,t}$, $m=1,...,M$, are either exogenous (nonrandom), or they follow the threshold (2.3) or the logistic process (2.4) (with $M=2$ in the latter),
@@
    $(\phi_{m},\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))\neq(\phi_{n},\text{vec}(A_{n,1}),...,\text{vec}(A_{n,p}))$ for all $m\neq n\in\{1,...,M\}$, and
  3. (3)
    there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
Then, for models with exogenous weights, the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$. For threshold models (2.3), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{m-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$. For logistic models (2.4), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,2$, and $c,\gamma$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ defined in Section 2 is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
@@
We have concluded that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1. In other words, at each $t$‚ changing the ordering or signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to an observationally distinct model. The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1. Identification of the AR and weight function parameters, in turn, is achieved only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$), because the identification fails in some special cases that comprise a set with Lebesgue measure zero. Condition (3) of Lemma 2 ensures that the transition weights exhibit enough variation to distinguish between different parameter values. If there are only two regimes, this condition is satisfied if transition weights take at least two distinct values.
If the impact matrix is time-invariant as in Lanne et al. (2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows from Lemma 2 that the structural shocks are identified up to ordering and signs. Such statistically identified shocks can then often be labelled as economic shocks based on external information (see Lanne et al., 2017). However, when the impact matrix varies in time, two complications arise. First, as the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified. Second, since the identification of $B_{y,t}$ at each $t$ is only up to ordering and signs of its columns, its unique identification requires that the parameters in the functional form of $B_{y,t}$ are restricted to fix the ordering and signs of the columns of $B_{y,t}$.
To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$. To fix the ordering and signs of the columns of $B_{y,t}$, such constraints need to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{1},...,B_{M}$ are observationally distinct at some $t$. However, constraints that are not restrictive in some specifications may be overidentifying in others, while constraints sufficient in some specifications may not be identifying in others, depending on the transition weights and shock distributions. Therefore, it is useful to divide the discussion on identification into the cases of binary (i.e., $\alpha_{m,t}\in\{0,1\}$) and more general transition weights.